OpenStudy (anonymous):

b) Determine the radius of convergence and the convergence interval (with analysis of the endpoints) for the power series \(\Large \sum_{n=0}^{\infty} \frac{x^{n}}{2^{(n^{2})}}\)

OpenStudy (anonymous):

\[ \frac{x^{n+1}}{\frac{2^{(n+1)^ 2} x^n}{2^{n^2}}}=\frac{x}{2^{ 2 n+1}} \]

OpenStudy (anonymous):

So R =Infinity

OpenStudy (anonymous):

\[ \lim_{n\to \infty} \frac{x^{n+1}}{\frac{2^{(n+1)^ 2} x^n}{2^{n^2}}}=\lim_{n\to \infty}\frac{x}{2^{ 2 n+1}}=0 \] for any x.

OpenStudy (anonymous):

is it our end solution mr Elias, i need to give my last homework tomorrow and need at least 6 points :)

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

ok thank you very much :)

OpenStudy (anonymous):

yw